Circular No. 8609

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 S4 (McNAUGHT) R. H. McNaught, Australian National University, reports his discovery

of a comet on CCD images taken with the 0.5-m Uppsala Schmidt telescope in the course of the Siding Spring Survey (discovery observation tabulated below); 120-s exposures taken on Oct. 1.6 UT clearly show a tail ∼ 8′′ long in p.a. ∼ 135^{o} (the discovery images on Sept. 30.6 show the comet as slightly diffuse).

2005 UT  α_{2000}  δ_{2000}  Mag.
Sept.30.55673  22^{h}06^{m}03.^{s}85  −37^{o}48′49′′.1  18.9

The available astrometry (including prediscovery Uppsala Schmidt images back to July 27), the following parabolic orbital elements, and an ephemeris appear on MPEC 2005-T12.

T = 2007 July 18.0603 TT  ω = 31.^{o}3980
Ω = 318.2820 }2000.0
q = 5.851768 AU  i = 107.9924

(3982) KASTEL P. Pravec and P. Kušnirák, Ondřejov Observatory; L. Kornoš and J.

Világi, Modra Observatory; D. Pray, Coventry, RI; R. Durkee, Minneapolis, MN; and W. Cooney, J. Gross and D. Terrell, Sonoita Research Observatory, AZ, report that photometric observations obtained during Sept. 24–29 reveal that (3982) has a lightcurve consisting of two linearly additive components with periods 8.488 and 5.835 (or possibly 2.918) hr and amplitudes 0.27 and 0.08 mag, respectively. No attenuations due to occultations/eclipses were seen, so the proposed interpretation of binary nature of the minor planet needs to be confirmed with further observations.

COMET C/2005 O2 (CHRISTENSEN)

Improved parabolic orbital elements from MPEC 2005-S57:

T = 2005 Sept. 8.4476 TT  ω = 263.^{o}8330
e = 0.859426  Ω = 280.7698 }2000.0
q = 3.333580 AU  i = 148.8919

a = 23.714057 AU n^{o} = 0.0085348 P = 115.5 years

2005 October 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green